Determine the electric potential at a point P which is 2.00 m from a point charge of -5.00µC.
1
Expert's answer
2020-06-04T09:56:24-0400

Given data

Charge is "q=-5.00*10^{-6} C"

The electric potential at a point P which is 2.00 m from a point charge of -5.00µC is

"V=\frac{kq}{r}=\frac{(9*10^9)(-5.00*10^{-6})}{2.00 }"

"=-2.25*10^4 V"